OpenStudy (anonymous):

Hey, I'm reviewing for my midterm, and I'm wondering how I can solve 2 cos^2 2x = 1 over all real numbers and interval (0, 2pi]?

myininaya (myininaya):

\[2 \cos^2(2x) =1 => \cos^2(2x)=\frac{1}{2}\] \[=> \cos(2x)=\pm \frac{1}{\sqrt{2}} => \cos(2x)=\pm \frac{\sqrt{2}}{2}\]

myininaya (myininaya):

now do you know when\[\cos(u)=\frac{\sqrt{2}}{2} \text{ or } \cos(u)=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

Well it could be both, because its in the interval (0, 2pi]?

OpenStudy (anonymous):

Oh wait I think I see what you're trying to ask.

myininaya (myininaya):

\[\cos(u)=\frac{\sqrt{2}}{2} \text{ when } u=\frac{\pi}{4}, \frac{5\pi}{4}\] \[\cos(u)=\frac{-\sqrt{2}}{2} \text{ when } u=\frac{3\pi}{4},\frac{7\pi}{4}\]

OpenStudy (anonymous):

Oh I see, thanks!

myininaya (myininaya):

so since u=2x, then you can solve both of those equations I just wrote

myininaya (myininaya):

for x

OpenStudy (anonymous):

So the denominators become 8?

myininaya (myininaya):

yep

myininaya (myininaya):

so you understood everything?

OpenStudy (anonymous):

yes! thank you <3

myininaya (myininaya):

:)